A flower bed is in the shape of a triangle with one side twice the length of the shortest side, and the third side is 19 feet more than the length of the shortest side. Find the dimensions if the perimeter is 159 feet.
A equals= ____ft (shortest sides
B equals= -ft (second side)
C equals= _____ ft (third side)

Answer by addingup(3677)

Shortest side:A
Other side:B and B= 2A
Third side:C and C= A+19
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
A+B+C = 159 Substitute for B and C
A+2A+A+19 = 159
4A = 140
A = 35 this is your shortest side. Then:
B = 35(2) = 70
c = 35+19 = 54